5. The regression line for the given data is hat(y) = -0.206x + 2.097.
( (x )(y) )
( ( -5 )(11) )
( ( -3 )(-6) )
( ( 4 )(8) )
( ( 1 )(-3) )
( ( -1 )(-2) )
( ( -2 )(1) )
( ( 0 )(5) )
( ( 2 )(-5) )
( ( 3 )(6) )
( ( -4 )(7) )
Determine the residual of a data point for which x = -5 and y = 11. (Points : 2)
14.127
7.873
3.127
-4.831

Answer by jim_thompson5910(35256)

y-hat = -0.206x + 2.097

y-hat = -0.206(-5) + 2.097

y-hat = 1.03 + 2.097

y-hat = 3.127

-------------------------------------------------------

Residual:

R = y - y-hat

R = 11 - 3.127

R = 7.873

So the answer is choice B